# Self-supervised pre-training with diffusion model for few-shot landmark detection in x-ray images

### v1

@@ 5.1 Tuning the DDPM pre-training iterations
The number of training iterations during the pre-training of the DDPM is a fundamental hyperparameter of our proposed approach. To evaluate the robustness and select the number of training iterations, we perform an experiment running three times the entire pipeline, considering $4k$, $6k$, $8k$, and $10k$ training iterations for the DDPM pre-training step.
Figure 3 visualizes the landmark detection results obtained in the validation sets for the Chest, Cephalometric, and Hand datasets, in terms of MRE. Our findings suggest that the pre-training is quite robust, except for the $4k$ iterations where the model is likely still learning crucial data features (as shown by the high standard deviation). Excessive training iterations beyond $8k$ may result in overfitting, and a slightly degrading performance on the downstream task. Considering these results, we decided to select the optimal number of training iterations for DDPM pre-training for subsequent experiments, utilizing $6k$ iterations for the Chest dataset, $8k$ iterations for the Cephalometric dataset, and $8k$ iterations for the Hand dataset.
@@ 5.2 Downstream task performance evaluation
In this section, we assess the effectiveness of our DDPM self-supervised pre-training method, benchmarking it with supervised ImageNet pre-training, and self-supervised state-of-the-art methods including MoCoV3, SimCLRV2, and DINO, across different numbers of labeled training samples (up to $50$) in the Chest, the Cephalometric, and the Hand datasets. Tables 1, 2, and 3 report the average test results for three independent runs on the three datasets (a bar chart version of the tables is provided in the Supplementary Material). With the exception of our approach, which implements a DDPM as described in the original paper [29], all reported results utilize DenseNet161 as the encoder backbone. We opt for this backbone due to its performance in the landmark detection task across all the datasets tested, based on a hold-out validation comparison with VGG19 and ResNeXt50_32x4D. Our approach outperforms both ImageNet and the explored alternative SSL approaches for all the tested datasets and quantities of training images in the downstream task. These results are qualitatively supported by Figure 4, providing a visual representation (i.e., the prediction heatmaps) of the methods’ performance with $10labels$, for each of the datasets included in our work.
Performance gains are particularly pronounced in low-data regimes, crucial for medical imaging applications where labeled data is often scarce.
In the Chest dataset (Table 1), with just one labeled sample, our method achieves an MRE of $14.99$ px compared to $143.67$ px for ImageNet and $64.69$ px for the best alternative (DINO). This represents a $89.6$% reduction in MRE versus ImageNet and a $76.8$% reduction versus DINO. The SDR at $9$px shows even more improvements, with our method reaching $64.63$%, far surpassing ImageNet (00%) and DINO ($23.44$%).
Similar trends are observed in the Cephalometric dataset (Table 2). Using one labeled sample, our method attains an MRE of $6.87$ mm versus $48.69$ mm for ImageNet and $24.39$ mm for the best alternative (MoCoV3) - reductions of $85.9$% and $71.8$%, respectively. The SDR at the restrictive 2mm further highlights our method’s superiority, achieving $49.43$% compared to ImageNet’s $0.12$% and MoCoV3’s $18.36$%.
The Hand dataset (Table 3) results reinforce these findings. With one labeled sample, our approach yields an MRE of $28.75$ mm versus $79.32$ mm for ImageNet and $97.31$ mm for the best alternative (DINO) - reductions of $63.8$% and $70.5$%, respectively. The SDR at 2mm shows our method at $27.87$%, significantly outperforming both ImageNet ($0.04$%) and DINO ($0.47$%).
As the number of labeled samples increases beyond $10$, the performance gap narrows, though our approach still maintains an edge. For instance, in the Cephalometric dataset with $25$ labeled samples, our method’s MRE of $1.62$ mm is closer to, but still better than, ImageNet’s $4.18$ mm and SimCLRV2’s $2.19$ mm. This trend is consistent across all datasets, indicating that while our method provides substantial benefits in extremely low-data scenarios, for which it is conceived and proposed, it continues to offer improvements even as more labeled data becomes available.
The obtained results highlight the efficacy of our DDPM self-supervised pre-training approach for landmark detection in few-shot learning scenarios, prevalent in medical imaging contexts where available high-quality annotated images are typically limited.
Finally, to frame our results in the state-of-the-art, we can consider that the current best-performing method [9] for the landmark detection task of the whole Chest, Cephalometric, and Hand datasets, consists of fine-tuning an ImageNet pre-trained model on the target datasets.
@@
Our method $\mathbf{14.99\pm 0.00}$ $\mathbf{19.92\pm 0.00}$ $\mathbf{46.34\pm 0.00}$ $\mathbf{64.63\pm 0.00}$
@@
Our method $\mathbf{6.04\pm 0.00}$ $\mathbf{32.52\pm 0.00}$ $\mathbf{65.04\pm 0.00}$ $\mathbf{86.99\pm 0.00}$
@@
Our method $\mathbf{5.33\pm 0.00}$ $\mathbf{33.74\pm 0.00}$ $\mathbf{67.89\pm 0.00}$ $\mathbf{86.59\pm 0.00}$
@@
Our method $\mathbf{4.62\pm 0.00}$ $\mathbf{32.93\pm 0.00}$ $\mathbf{73.17\pm 0.00}$ $\mathbf{91.87\pm 0.00}$
@@
Our method $\mathbf{4.31\pm 0.01}$ $\mathbf{39.63\pm 0.61}$ $\mathbf{80.08\pm 0.41}$ $\mathbf{93.50\pm 0.41}$
@@
$1$ ImageNet $48.69\pm 0.03$ $0.12\pm 0.01$ $0.16\pm 0.01$ $0.21\pm 0.02$ $0.41\pm 0.02$
MoCoV3 $24.39\pm 0.75$ $18.36\pm 0.35$ $23.92\pm 0.84$ $28.55\pm 0.93$ $35.89\pm 1.11$
SimCLRV2 $49.27\pm 9.00$ $4.14\pm 2.01$ $5.42\pm 2.56$ $6.73\pm 3.04$ $9.13\pm 4.10$
DINO $30.78\pm 0.73$ $7.64\pm 0.38$ $10.45\pm 0.50$ $13.57\pm 0.52$ $18.68\pm 0.64$
Our method $\mathbf{6.87\pm 0.00}$ $\mathbf{49.43\pm 0.00}$ $\mathbf{58.19\pm 0.00}$ $\mathbf{66.32\pm 0.00}$ $\mathbf{74.78\pm 0.00}$
$5$ ImageNet $46.84\pm 0.03$ $0.14\pm 0.02$ $0.22\pm 0.01$ $0.29\pm 0.00$ $0.52\pm 0.01$
MoCoV3 $7.53\pm 0.20$ $46.17\pm 0.41$ $55.26\pm 0.64$ $64.29\pm 0.53$ $75.48\pm 0.40$
SimCLRV2 $13.73\pm 0.16$ $33.69\pm 0.18$ $41.48\pm 0.35$ $49.31\pm 0.31$ $59.40\pm 0.10$
DINO $12.85\pm 0.22$ $35.88\pm 0.09$ $43.80\pm 0.27$ $50.65\pm 0.15$ $60.20\pm 0.32$
Our method $\mathbf{2.69\pm 0.00}$ $\mathbf{64.72\pm 0.00}$ $\mathbf{75.16\pm 0.00}$ $\mathbf{83.18\pm 0.00}$ $\mathbf{90.59\pm 0.00}$
$10$ ImageNet $27.44\pm 0.23$ $10.18\pm 0.03$ $13.33\pm 0.25$ $16.96\pm 0.20$ $22.68\pm 0.35$
MoCoV3 $5.32\pm 0.02$ $56.22\pm 0.29$ $66.91\pm 0.29$ $75.81\pm 0.13$ $85.25\pm 0.06$
SimCLRV2 $10.09\pm 0.07$ $43.67\pm 0.15$ $52.84\pm 0.09$ $61.19\pm 0.03$ $71.27\pm 0.20$
DINO $8.73\pm 0.08$ $48.19\pm 0.15$ $57.70\pm 0.27$ $65.90\pm 0.15$ $75.64\pm 0.00$
Our method $\mathbf{1.93\pm 0.00}$ $\mathbf{72.25\pm 0.00}$ $\mathbf{82.48\pm 0.00}$ $\mathbf{89.83\pm 0.00}$ $\mathbf{96.13\pm 0.00}$
$25$ ImageNet $4.18\pm 0.01$ $57.97\pm 0.37$ $69.03\pm 0.46$ $78.04\pm 0.20$ $87.11\pm 0.09$
MoCoV3 $2.22\pm 0.01$ $73.21\pm 0.20$ $82.69\pm 0.07$ $89.98\pm 0.25$ $95.79\pm 0.08$
SimCLRV2 $2.19\pm 0.01$ $69.35\pm 0.21$ $80.40\pm 0.04$ $88.36\pm 0.15$ $95.39\pm 0.17$
DINO $3.02\pm 0.06$ $68.22\pm 0.33$ $78.70\pm 0.27$ $86.81\pm 0.07$ $93.46\pm 0.09$
Our method $\mathbf{1.62\pm 0.03}$ $\mathbf{77.33\pm 1.12}$ $\mathbf{85.84\pm 1.49}$ $\mathbf{91.71\pm 1.20}$ $\mathbf{96.86\pm 0.63}$
$50$ ImageNet $1.60\pm 0.04$ $81.17\pm 0.42$ $89.20\pm 0.43$ $94.49\pm 0.32$ $98.29\pm 0.12$
MoCoV3 $1.63\pm 0.01$ $81.72\pm 0.04$ $89.72\pm 0.08$ $94.76\pm 0.02$ $98.28\pm 0.05$
SimCLRV2 $1.67\pm 0.02$ $79.55\pm 0.24$ $87.65\pm 0.27$ $93.67\pm 0.03$ $97.84\pm 0.01$
DINO $1.65\pm 0.01$ $80.89\pm 0.02$ $88.47\pm 0.48$ $93.75\pm 0.06$ $97.82\pm 0.07$
Our method $\mathbf{1.40\pm 0.01}$ $\mathbf{82.16\pm 1.16}$ $\mathbf{89.98\pm 0.62}$ $\mathbf{94.63\pm 0.54}$ $\mathbf{98.30\pm 0.09}$
@@
Our method $\mathbf{1.74\pm 0.03}$ $\mathbf{78.04\pm 0.94}$ $\mathbf{96.35\pm 0.51}$ $\mathbf{99.02\pm 0.04}$
Figure 4: Comparison of landmark detection performance between our pre-training method and other self-supervised and ImageNet pre-training approaches on Chest, Cephalometric, and Hand x-ray test sets, using only 10 labeled training samples.
Table 4: Landmark detection test performance of our method, pre-trained on the Hand dataset and fine-tuned on the Chest / Cephalometric datasets, across varying numbers of labeled training images.
@@ 5.3 Impact of different pre-training datasets
In this section, we explore whether a different in-domain dataset can still boost the performance in the potential situation where both the number of available images and annotations are limited. To this purpose, we select the best DDPM model pre-trained on the larger dataset (Hand), and we fine-tune it on the other two smaller datasets (Chest and Cephalometric), with the same protocol as the previous experiments (three runs and different numbers of annotated training images for the downstream task). Table 4 reports the obtained results. Our experiments reveal the efficacy of using a larger in-domain dataset (Hand) for pre-training in scenarios with limited target task data.
Using a model pre-trained on the Hand dataset for the Chest landmark detection task yields better results compared to a model pre-trained on the Chest dataset itself for most training samples quantity scenarios (5 to 50 labeled images). Even with just one labeled sample, the Hand pre-trained model achieves comparable performance to the Chest pre-trained model.
For the Cephalometric landmark detection task, the Hand pre-trained model does not outperform the Cephalometric pre-trained model, but it does achieve competitive results across all the selected number of training samples. The performance gap between the two pre-training approaches is relatively small, indicating that the Hand pre-training still transfers useful features to the Cephalometric task. Importantly, our approach consistently outperforms other pre-training methods (ImageNet, MoCoV3, SimCLRV2, DINO), regardless of the number of labeled training samples, for both tasks. This underscores the value of in-domain pre-training using our approach, even when the pre-training dataset differs from the target task. Our goal is to utilize images from the same dataset as the downstream task, thereby simulating real-world scenarios with limited data and label availability. Furthermore, we focused on label-efficient pre-training while minimizing computational resource usage, given that training diffusion models on extensive unlabeled datasets demands substantial resources and time. Even though different larger unlabeled datasets might enhance pre-training outcomes, our main point is that using a limited unannotated dataset can still surpass self-supervised and supervised state-of-the-art pre-training techniques like ImageNet (which has one million images), when images and annotations availability are constrained.

### v3

@@ 5.1 Tuning the DDPM pre-training iterations
The number of training iterations during DDPM pre-training is a fundamental hyperparameter in our approach. To evaluate its impact and determine the optimal value, we conducted an experiment running the entire pipeline three times, considering $4k$, $6k$, $8k$, and $10k$ iterations for the DDPM pre-training step. Figure 2 illustrates the landmark detection results obtained on the validation sets for the Chest, Cephalometric, and Hand datasets in terms of MRE.
Our findings indicate that the pre-training is quite robust, except for the $4k$ iterations where the model is likely still learning crucial data features (as shown by the high standard deviation). Excessive training iterations beyond $8k$ may result in overfitting, resulting in slightly degraded performance on the downstream task. Based on these results, we selected the optimal number of training iterations for DDPM pre-training in subsequent experiments, utilizing $6k$, $8k$, and $8k$ iterations for the Chest, Cephalometric, and Hand datasets respectively.
@@ 5.2 Downstream task performance evaluation
We assess the effectiveness of our DDPM self-supervised pre-training method by benchmarking it against supervised ImageNet pre-training and self-supervised state-of-the-art methods (MoCoV3, SimCLRV2, and DINO) across different numbers of labeled training samples (1, 5, 10, 25, and 50) in the Chest, Cephalometric, and Hand datasets. Additionally, we explore an ablation of our method using the contour-hugging approach [20] as an alternative framework for the downstream task. Tables 1, 2, and 3 report the average test results for three independent runs (a bar chart version of the tables is provided in the Supplementary Material). Except for our approach, which implements a DDPM as described in [12], all methods use DenseNet161 as the encoder backbone, chosen for its performance across all datasets based on hold-out validation comparisons with VGG19 and ResNeXt50_32x4D.
Our approach consistently outperforms both ImageNet and alternative SSL approaches across all datasets and training image quantities. Figure 3 provides visual support, showing prediction heatmaps for each method with 10 labels across all datasets. Performance gains are particularly pronounced in low-data regimes. For instance, in the Chest dataset (Table 1) with one labeled sample, our method achieves an MRE of $14.99$px compared to $143.67$px for ImageNet and $64.69$px for DINO (the best alternative), representing reductions of $89.6$% and $76.8$% respectively. The SDR at $9$px shows even greater improvements, with our method reaching $64.63$%, far surpassing other approaches.
@@
Our method with [20] $50.03\pm 0.00$ $\mathbf{21.14\pm 0.00}$ $43.90\pm 0.00$ $50.81\pm 0.00$
Our method $\mathbf{14.99\pm 0.00}$ $19.92\pm 0.00$ $\mathbf{46.34\pm 0.00}$ $\mathbf{64.63\pm 0.00}$
@@
Our method with [20] $8.14\pm 0.00$ $\mathbf{60.16\pm 0.00}$ $\mathbf{82.11\pm 0.00}$ $\mathbf{86.99\pm 0.00}$
Our method $\mathbf{6.04\pm 0.00}$ $32.52\pm 0.00$ $65.04\pm 0.00$ $\mathbf{86.99\pm 0.00}$
@@
Our method with [20] $\mathbf{4.47\pm 0.00}$ $\mathbf{60.98\pm 0.00}$ $\mathbf{88.62\pm 0.00}$ $\mathbf{93.50\pm 0.00}$
Our method $5.33\pm 0.00$ $33.74\pm 0.00$ $67.89\pm 0.00$ $86.59\pm 0.00$
@@
Our method with [20] $\mathbf{2.92\pm 0.00}$ $\mathbf{76.02\pm 0.00}$ $\mathbf{94.31\pm 0.00}$ $\mathbf{96.34\pm 0.00}$
Our method $4.62\pm 0.00$ $32.93\pm 0.00$ $73.17\pm 0.00$ $91.87\pm 0.00$
@@
Our method with [20] $\mathbf{2.69\pm 0.00}$ $\mathbf{76.83\pm 0.00}$ $\mathbf{93.50\pm 0.00}$ $\mathbf{97.56\pm 0.00}$
Our method $4.31\pm 0.01$ $39.63\pm 0.61$ $80.08\pm 0.41$ $93.50\pm 0.41$
Similar trends are observed in the Cephalometric (Table 2) and Hand datasets (Table 3). With one labeled sample in the Cephalometric dataset, our method attains an MRE of $15.71$mm versus $86.71$mm for ImageNet and $43.49$mm for MoCoV3, with corresponding improvements in SDR. In the Hand dataset, our approach yields an MRE of $28.75$mm versus $79.32$mm for ImageNet and $97.31$mm for DINO, again with SDR improvements. As the number of labeled samples increases, the performance gap narrows, though our approach maintains an edge. For example, in the Cephalometric dataset with $25$ labeled samples, our method’s MRE of $2.84$mm still outperforms ImageNet’s $7.67$mm and SimCLRV2’s $3.99$mm. This trend is consistent across all datasets, indicating that our method provides substantial benefits in extremely low-data scenarios while continuing to offer improvements as more labeled data becomes available. These results highlight our DDPM self-supervised pre-training’s effectiveness for landmark detection in few-shot learning scenarios, common in medical imaging where high-quality annotated images are limited.
@@
$1$ ImageNet $86.71\pm 0.02$ $0.02\pm 0.00$ $0.06\pm 0.00$ $0.06\pm 0.00$ $0.15\pm 0.00$
MoCoV3 $43.49\pm 1.21$ $6.79\pm 0.10$ $10.83\pm 0.25$ $13.67\pm 0.30$ $19.99\pm 0.30$
SimCLRV2 $100.68\pm 12.18$ $0.82\pm 0.42$ $1.27\pm 0.68$ $1.63\pm 0.79$ $2.51\pm 1.19$
DINO $58.67\pm 1.22$ $2.47\pm 0.14$ $4.42\pm 0.07$ $5.52\pm 0.07$ $8.60\pm 0.30$
Our method with [20] $20.13\pm 0.00$ $\mathbf{18.06\pm 0.00}$ $25.14\pm 0.00$ $31.94\pm 0.00$ $41.37\pm 0.00$
Our method $\mathbf{15.71\pm 0.00}$ $17.31\pm 0.00$ $\mathbf{27.14\pm 0.00}$ $\mathbf{33.24\pm 0.00}$ $\mathbf{45.14\pm 0.00}$
$5$ ImageNet $84.40\pm 0.02$ $0.06\pm 0.00$ $0.11\pm 0.00$ $0.15\pm 0.00$ $0.17\pm 0.00$
MoCoV3 $13.44\pm 0.12$ $18.62\pm 0.19$ $29.56\pm 0.18$ $35.99\pm 0.02$ $49.38\pm 0.16$
SimCLRV2 $24.59\pm 0.12$ $13.06\pm 0.31$ $21.24\pm 0.25$ $26.44\pm 0.23$ $36.52\pm 0.37$
DINO $23.83\pm 0.21$ $14.00\pm 0.26$ $23.01\pm 0.54$ $28.42\pm 0.52$ $39.13\pm 0.38$
Our method with [20] $11.53\pm 0.00$ $\mathbf{45.05\pm 0.00}$ $\mathbf{54.00\pm 0.00}$ $\mathbf{61.03\pm 0.00}$ $\mathbf{70.55\pm 0.00}$
Our method $\mathbf{8.30\pm 0.00}$ $27.18\pm 0.00$ $42.65\pm 0.00$ $51.01\pm 0.00$ $66.34\pm 0.00$
$10$ ImageNet $50.09\pm 0.44$ $3.18\pm 0.16$ $5.92\pm 0.25$ $7.49\pm 0.26$ $11.19\pm 0.42$
MoCoV3 $9.30\pm 0.03$ $23.84\pm 0.12$ $37.20\pm 0.13$ $45.34\pm 0.17$ $60.18\pm 0.13$
SimCLRV2 $18.54\pm 0.09$ $17.80\pm 0.29$ $28.16\pm 0.04$ $34.65\pm 0.22$ $46.74\pm 0.23$
DINO $15.64\pm 0.26$ $18.86\pm 0.12$ $30.89\pm 0.10$ $38.22\pm 0.34$ $51.83\pm 0.45$
Our method with [20] $5.21\pm 0.00$ $\mathbf{52.11\pm 0.00}$ $\mathbf{62.97\pm 0.00}$ $\mathbf{71.43\pm 0.00}$ $\mathbf{81.68\pm 0.00}$
Our method $\mathbf{3.52\pm 0.00}$ $32.06\pm 0.00$ $48.59\pm 0.00$ $58.97\pm 0.00$ $74.69\pm 0.00$
$25$ ImageNet $7.67\pm 0.39$ $24.25\pm 0.10$ $37.83\pm 0.13$ $45.58\pm 0.26$ $62.56\pm 0.35$
MoCoV3 $4.04\pm 0.03$ $34.81\pm 0.18$ $51.26\pm 0.14$ $60.80\pm 0.03$ $76.78\pm 0.10$
SimCLRV2 $3.99\pm 0.04$ $30.63\pm 0.23$ $47.04\pm 0.18$ $56.20\pm 0.13$ $73.29\pm 0.21$
DINO $5.38\pm 0.19$ $30.00\pm 0.20$ $45.98\pm 0.04$ $55.36\pm 0.06$ $72.30\pm 0.19$
Our method with [20] $\mathbf{2.72\pm 0.00}$ $\mathbf{65.47\pm 0.00}$ $\mathbf{75.77\pm 0.00}$ $\mathbf{82.55\pm 0.00}$ $\mathbf{90.61\pm 0.00}$
Our method $2.84\pm 0.00$ $39.84\pm 0.01$ $57.61\pm 0.01$ $67.17\pm 0.01$ $81.01\pm 0.00$
$50$ ImageNet $2.82\pm 0.09$ $42.93\pm 2.66$ $60.74\pm 3.48$ $69.69\pm 3.28$ $83.87\pm 2.24$
MoCoV3 $2.98\pm 0.01$ $43.68\pm 0.32$ $61.84\pm 0.13$ $70.48\pm 0.03$ $84.81\pm 0.19$
SimCLRV2 $3.17\pm 0.04$ $39.94\pm 0.19$ $58.70\pm 0.08$ $67.66\pm 0.05$ $81.85\pm 0.10$
DINO $2.98\pm 0.02$ $40.33\pm 0.38$ $59.06\pm 0.39$ $68.99\pm 0.42$ $83.78\pm 0.64$
Our method with [20] $\mathbf{2.13\pm 0.00}$ $\mathbf{73.60\pm 0.00}$ $\mathbf{81.22\pm 0.00}$ $\mathbf{86.78\pm 0.00}$ $\mathbf{92.70\pm 0.00}$
Our method $2.50\pm 0.00$ $44.91\pm 0.05$ $62.84\pm 0.02$ $72.82\pm 0.00$ $85.81\pm 0.02$
@@
Our method with [20] $7.16\pm 0.00$ $\mathbf{79.38\pm 0.00}$ $88.36\pm 0.00$ $90.19\pm 0.00$
Our method $\mathbf{1.74\pm 0.03}$ $78.04\pm 0.94$ $\mathbf{96.35\pm 0.51}$ $\mathbf{99.02\pm 0.04}$
Figure 3: Comparison of landmark detection performance between our DDPM pre-training method and alternative self-supervised and ImageNet pre-training approaches on Chest, Cephalometric, and Hand x-ray test sets, using only 10 labeled training samples.
Table 4: Comparison of our method vs. state-of-the-art YOLO framework using 1, 5, and 10 labeled samples. + indicates mixed dataset training. Both methods utilizes input images resized at $256$x256256256256 pixels. Our method performs competitively using only 1 labeled sample.
Methods Number of Labeled Training Samples Chest Cephalometric Hand
MRE $\downarrow$ SDR(%) $\uparrow$ MRE $\downarrow$ SDR(%) $\uparrow$ MRE $\downarrow$ SDR(%) $\uparrow$
(px) 3px 6px 9px (mm) 2mm 2.5mm 3mm 4mm (mm) 2mm 4mm 10mm
YOLO+ 10 22.55 32.11 63.82 76.42 15.05 18.23 26.50 39.62 53.49 52.43 11.59 19.52 24.14
YOLO+ 5 23.13 23.57 50.40 67.07 19.38 11.53 17.07 25.95 35.53 64.24 15.15 24.91 31.79
YOLO+ 1 84.37 4.07 13.41 19.51 53.80 7.35 11.22 17.26 23.39 61.36 8.26 17.16 31.47
Our method 1 14.99 19.92 46.34 64.63 15.71 17.31 27.14 33.24 45.14 28.75 27.87 46.44 58.75
Table 5: Landmark detection test performance of our Gaussian-heatmaps method, pre-trained on the Hand dataset and fine-tuned on the Chest / Cephalometric datasets, across varying numbers of labeled training images.
@@
Finally, to contextualize our results, we compare our method to YOLO [44], a state-of-the-art universal anatomical landmark detection model. YOLO utilizes mixed dataset training and combines a universal U-Net for local feature learning with a global network for landmark disambiguation. As Table 4 shows, our method performs comparably or better than YOLO variants across all three datasets. Despite YOLO’s mixed dataset training (denoted by ’$+$’), our approach achieves competitive results with just one labeled sample, even with $5$ and $10$ samples.
Ablation study on the downstream task setting.
We perform an ablation study comparing the performance of our downstream task framework, with the one introduced in [20] (second-last entry in our results Tables), mainly differing in the usage of a single annotated pixel to generate unrestricted contour-hugging heatmaps and a negative log-likelihood loss. For the Hand dataset, which had more extensive pre-training data, our approach consistently outperforms the evaluated alternative. However, the contour-hugging method generally shows better SDR performance on the Chest and the Cephalometric datasets. Regarding MRE, our original framework performs better up to $10$ training samples, showing lower performance with more annotated data. These findings suggest that the effectiveness of each downstream setting depends on pre-training data volume and quality, as well as dataset-specific characteristics.
@@ 5.3 Impact of different pre-training datasets
In this section, we investigate whether a different in-domain dataset can enhance performance when both available images and annotations are limited. We select the best DDPM model pre-trained on the larger Hand dataset and fine-tune it on the smaller Chest and Cephalometric datasets, following the same protocol as previous experiments with Gaussian heatmaps. Table 5 reports the results.
Our experiments show the effectiveness of using a larger in-domain dataset (Hand) for pre-training in scenarios with limited target task data. For the Chest landmark detection task, the Hand pre-trained model outperforms the Chest pre-trained model in most scenarios (5 to 50 labeled images). Even with just one labeled sample, it achieves comparable performance to the Chest pre-trained model. For the Cephalometric landmark detection task, while the Hand pre-trained model doesn’t surpass the Cephalometric pre-trained model, it achieves competitive results across all training sample quantities. The small performance gap indicates that Hand pre-training still transfers useful features to the Cephalometric task. Notably, our approach consistently outperforms other pre-training methods (ImageNet, MoCoV3, SimCLRV2, DINO) and frameworks (YOLO) for both tasks, regardless of the number of labeled training samples. This underscores the value of in-domain pre-training using our approach, even when the pre-training dataset differs from the target task.
Our goal is to simulate real-world scenarios with limited data and label availability by utilizing images from the same dataset as the downstream task. We focused on label-efficient pre-training while minimizing computational resource usage, given the substantial resources and time required for training diffusion models on extensive unlabeled datasets. While larger unlabeled datasets might enhance pre-training outcomes, our key finding is that using a limited unannotated dataset can still surpass self-supervised and supervised state-of-the-art pre-training techniques like ImageNet (which has one million images) when image and annotation availability is constrained.